OpenStudy (anonymous):

Simplify this expression and state any restrictions on the variables. (x+1/x^2+2x-8) - (x/4x-8) please show me step by step!

OpenStudy (e.mccormick):

\[\frac{x+1}{x^2+2x-8} - \frac{x}{4x-8}\] OK, where are you getting stuck?

OpenStudy (anonymous):

\[\frac{ x+1 }{ x^{2}+2x-8 }-\frac{ x }{ 4x-8}\]

OpenStudy (anonymous):

I have no idea how to do these!!

OpenStudy (e.mccormick):

Start by factoring the denominators. you need to find a comon denominator and the only way to do so is with factors.

OpenStudy (anonymous):

\[\frac{ x + 1 }{ (x -2)(x-4) } - \frac{ x }{ 4(x-2) }\]

OpenStudy (e.mccormick):

Ta da... factors. Hehe. You see what superdav815 has done there?

OpenStudy (anonymous):

yeah but the fraction part confuses me, i stink at fractions :(

OpenStudy (anonymous):

\[\frac{ 4(x+1) }{ 4(x-2)(x+4) } - \frac{ x(x+4) }{ 4(x-2)(x+4) }\]

OpenStudy (anonymous):

you can do the rest

OpenStudy (e.mccormick):

Well, it is the same old common deniminator thing. Just like when you want to add 1/2 and 1/3.

OpenStudy (e.mccormick):

What you look for in the denominator of each fraction is overlap, what is there and what is missing. That is how superdav815 got what he did. Do you understamd that, or would you like me to show it to you a different way?

OpenStudy (anonymous):

like i know how to solve 4(x+1) and problems like that by itself but i don't get hoe to do it when you start putting them on top of each other lol

OpenStudy (e.mccormick):

No problem at all! Then you know the foundation of it,, you jsut need some specifics. Are you OK with the factoring part?

OpenStudy (anonymous):

factoring is when you solve 4(x+1) thats called factoring right?

OpenStudy (e.mccormick):

No, that is called solving. Facroing is when you turn \(x^2+2x-8\) into \((x−2)(x−4)\). Another name for it could be un-multiplying.

OpenStudy (anonymous):

oh lol sorry yeah i understand how to do that, i just forget how to start sometimes

OpenStudy (e.mccormick):

NP. The next part is one of the ones that might confuse you, lowest comon denominator. I am doing something to deal with that, but have you heard of it and are you at least somewhat familliar with the idea?

OpenStudy (anonymous):

yeah

OpenStudy (e.mccormick):

Good, good. That process is the key to these. Almost done with my drawing. Then we can talk about any parts you do not get.

OpenStudy (anonymous):

okay

OpenStudy (e.mccormick):

OK. Here is is!

OpenStudy (e.mccormick):

That is the foundation behind what happend that let superdav815 just post those equations. He found the lowest common denominator, or LCD, then looked for what each side was missing and multiplied by one to add it in. Now the one he used is things like \(\frac{4}{4}\), but that is still one.

OpenStudy (anonymous):

yeah i understand that part now, so what do I do next?

OpenStudy (e.mccormick):

Subtract.

OpenStudy (e.mccormick):

What does the top become?

OpenStudy (anonymous):

4x+5-x^2+4x?

OpenStudy (e.mccormick):

I think you distributed things wrong.

OpenStudy (anonymous):

darn -_- what did i do wrong?

OpenStudy (e.mccormick):

Lets take it apart and try again. \(4(x+1)-[x(x+4)]\) So lets grab: \(4(x+1)\) What does that distribute to?

OpenStudy (anonymous):

4x+4

OpenStudy (e.mccormick):

Yes, good. Now: \(x(x+4)\)

OpenStudy (anonymous):

1x^2+4x

OpenStudy (e.mccormick):

Perfect. OK, if I put those back in I get: \(4x+4-[x^2+4x]\) But I still need to distribute the \(-1\) that is hiding outside the [.

OpenStudy (anonymous):

how do you do that again?

OpenStudy (e.mccormick):

Multiply the negative through. \(-[x^2+4x]\) So the negativd distributes.

OpenStudy (anonymous):

-1x^2+3x

OpenStudy (e.mccormick):

When you distribute a nagative, it changes the sign of each term inside. Not the value, just the sign.

OpenStudy (anonymous):

okay so it would be -1x^2-4x

OpenStudy (e.mccormick):

YES! OK, so, let me put that all back in the top, we are just looking at the top of the fraction and ignoring the bottom for now. \(4x+4-x^2-4x\) Now, see any like terms to combined?

OpenStudy (anonymous):

1x^2+4-x^2

OpenStudy (e.mccormick):

\(1x^2+4-x^2\) ? Not sure how that happned. I was taling about the \(4x\) and \(-4x\). The rest, well, is unique.

OpenStudy (e.mccormick):

It looks like you are having a little trouble with polynomials. I have done tons of them and still make mistakes at times, so I am not going to say it is surprising. I think we all have troubles with one part of math or another. However, I have a sneaking suspicion at this point it is just getting frustrating. My aim with this top of the fraction is this: \(4x+4-x^2-4x\implies 4x-4x+4-x^2 \implies 4-x^2\) Do you see what I did there?

OpenStudy (anonymous):

i think so

OpenStudy (e.mccormick):

It is just the 4x and -4x that cancel. OK. Let me put what is left back into the fraction. There is still a little work to do, but it is close to done.

OpenStudy (e.mccormick):

That gets us to: \[\frac{4-x^2}{4(x-2)(x+4)}\]

OpenStudy (e.mccormick):

Now, all that is left is the simplify part. All the subtracting is done.

OpenStudy (anonymous):

okay 4x-8+4x+16 -> 8x^2+22?

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

@superdav815 Hehe. What, not going to draw that in the machine like mine?

OpenStudy (anonymous):

srry

OpenStudy (e.mccormick):

@sophiasmommy94 Oh, you multiplied it out. Actually, the top factors a bit. Look at the pic SD uploaded. It has the end in there.

OpenStudy (e.mccormick):

@superdav815 Just teasing. Nice handwriting.

OpenStudy (anonymous):

lol thank you

OpenStudy (anonymous):

forgot to add in the picture: another restriction is \[x \neq -4\]

OpenStudy (e.mccormick):

Yah, from the original. Both the start and the end must be true.

OpenStudy (anonymous):

You two have been such a BIG help!! Thank you Thank you Thank you!!! Do any of you have time for one more question?

OpenStudy (anonymous):

whats the question

OpenStudy (e.mccormick):

I need to head to bed. But I did one to share one last thing on this one. What the restrictions mean. See, people hear it is invalid, or does not exist, but that does not always mean much. Here is a graph of the two equations that were subtracted: https://www.desmos.com/calculator/qldnssqthu Notice the big gaps where the graph suddenly shoots up and down at the same time? Those are the restriction points for the original graph. I could also add the answer, or you can, and see where it suddenly goes funky.

OpenStudy (anonymous):

simplify: \[\frac{ 1 }{ 1+\frac{ x }{ y } }\]

OpenStudy (anonymous):

this is what i got so far 1/(1+y/x)= 1/(x/x+y/x)= 1/(x+y/x)

OpenStudy (anonymous):

Thank you @e.mccormick

OpenStudy (e.mccormick):

np. Have fun!

OpenStudy (e.mccormick):

Watch your common denominator in that first step you did. It is a y they need to share, not an x.

OpenStudy (anonymous):

\[\frac{ 1 }{ 1 +\frac{ x }{ y } } = \frac{ 1 (y)}{( 1 +\frac{ x }{ y } )(y) }\]

OpenStudy (anonymous):

do you get that last step? I just multiplied numerator and denominator by y

OpenStudy (anonymous):

yeah, so it's simplified?

OpenStudy (anonymous):

im guessing the answer they want is this \[\frac{ y }{ y+x }\]

OpenStudy (anonymous):

okay, thank you!!

OpenStudy (anonymous):

can you write out the steps so i can put them in my notebook for future refrence?

OpenStudy (anonymous):

@superdav815

OpenStudy (anonymous):

thank you!!